# Coordination Mechanisms with Rank-Based Utilities

### v1

@@ Abstract
In traditional scheduling problems, a central authority assigns jobs to machines to optimize a global objective. In practice, scheduling environments are often managed by multiple strategic users, each seeking to maximize their individual benefit. Job-scheduling games model these decentralized environments, where players (jobs) select their own assignment, each aiming to optimize its own utility. Our work considers a model in which every machine is associated with a scheduling policy, which determines the order according to which it processes the jobs assigned to it. Such games, also denoted coordination mechanisms, were analyzed in the past assuming that the objective of every job is to minimize its completion time. Our work introduces and studies games with rank-based utilities, where players aim to perform well relative to their competitors rather than solely focusing on completion time.
We first show that with rank-based utilities, even a simple game on two identical machines and unit-length jobs, may not have a Nash equilibrium (NE) assignment. This motivates the study of the problem of determining, given a game instance, whether a pure NE exists. We show that this problem is NP-complete, even for very restricted classes. On the other hand we identify several classes of games for which a NE is guaranteed to exist, or the decision problem can be solved in polynomial time. Our analysis distinguishes between $(i)$ games played in an environment of identical vs. related parallel machines, $(ii)$ games played in an environment in which machines may have different scheduling policies vs. an environment with a global priority list - where all the machines share the same policy.
For games with unit-length jobs and a global priority list, we provide a clear characterization of instances that admit a NE. For games that have no NE, we propose a dynamic of best-response deviations with a fast convergence to an efficient sink equilibrium. Additionally, we show that for games with unit-length jobs and two related machines, deciding and constructing a NE can be done in linear time. We also introduce the Inversed-Policies coordination mechanism, which guarantees a NE in games with two identical machines and arbitrary job lengths.
Our analysis explores how competition influences the efficiency of equilibria, extending classical efficiency bounds from traditional scheduling games to the rank-based settings. We show that competition can either improve or worsen performance and show that the competition structure has a crucial influence on both the game stability, convergence of best-response dynamics, and the equilibrium inefficiency.
Our findings offer new insights into the existence and efficiency of equilibria in competitive job-scheduling environments.
@@ 1.3 Related Work
Scheduling games were initially studied in the setting in which each machine processes its jobs in parallel so that the completion time of each job is equal to the total load on the machine. The corresponding papers analyze the inefficiency of selfish behavior with respect to the makespan. Czumaj and Vöcking [11] gave tight bounds on the price of anarchy for related machines, whereas Awerbuch et al. [3] and Gairing et al. [15] provided tight bounds for restricted machine settings. These tight bounds grow with the number of machines and motivated the study of coordination mechanisms, i.e., local scheduling policies, to reduce the price of anarchy.
@@ 3.2 Related Machines
    $L_{1}(s)=L_{2}(s)$, and each of the two last jobs is prioritized on its machine over the other last job. Formally, $j_{1}\prec_{1}j_{2}$ and $j_{2}\prec_{2}j_{1}$.
@@
    $L_{1}(s)>L_{2}(s)$, and either no job on $M_{1}$ has completion time $C_{j_{2}}(s)$, or for the job $j^{\prime}$ such that $C_{j^{\prime}}(s)=C_{j_{2}}(s)$ it holds that $j^{\prime}\prec_{1}j_{2}$.
Proof:    We start by showing that each of these conditions ensures no job satisfies the conditions outlined in Claim 3.1. First, if the slow machine is more loaded, then clearly, $C_{j_{2}}(s)$ is not equal to the completion time of any job on $M_{1}$. Also, $C_{j_{1}}(s)$ is during the interval in which $j_{2}$ is processed on $M_{2}$. Second, if $L_{1}(s)=L_{2}(s)$ and $j_{1}\prec_{1}j_{2}$ and $j_{2}\prec_{2}j_{1}$, then none of the two jobs has a beneficial deviation, and no other job is last on its machine. Finally, if the third condition is valid, then as required, no last job has a same-rank job on the other machine, such that a deviation is beneficial.
@@ 5 Related Machines - Equilibrium Existence and Computation
If $0<c<\left\lceil\frac{b}{a}\right\rceil$, assume that Algorithm 2 is executed. After the first $\ell\cdot(a+b)-2$ jobs are assigned, we enter the $({\ell\cdot(a+b)-1})^{th}$ iteration. Let $j_{1}$ be the first unassigned job in $\pi_{1}$. In Step $3$ of the algorithm, machine $i^{*}$ might be either $M_{1}$ or $M_{2}$. Suppose the tiebreaker resolves in favor of assigning $j_{1}$ to $M_{1}$. Let $s$ be the produced schedule, and let $j_{2}$ be the last job processed on $M_{2}$. Since $0<c<\left\lceil\frac{b}{a}\right\rceil$, it follows that $L_{1}(s)>L_{2}(s)$. Note that $C_{j_{1}}(s)=C_{j_{2}}(s)$ and $j_{1}\prec_{1}j_{2}$. Therefore, by Theorem 3.3 $s$ is a NE.
@@ 5.1 Global Priority List
For the other direction, let $s$ be a NE schedule, and let $j_{1}$ and $j_{2}$ denote the last jobs processed on $M_{1}$ and $M_{2}$, respectively. Assume by contradiction that $c=0$. Schedule $s$ must be balanced, since otherwise, the last job can benefit from migration. Therefore, $L_{1}(s)=L_{2}(s)$, and by Theorem 3.3, since $\pi$ is global, $s$ is not a NE.

### v2

@@ Abstract
In classical job-scheduling games, each job behaves as a selfish player, choosing a machine to minimize its own completion time. To reduce the equilibria inefficiency, coordination mechanisms are employed, allowing each machine to follow its own scheduling policy. In this paper we study the effects of incorporating rank-based utilities within coordination mechanisms across environments with either identical or unrelated machines. With rank-based utilities, players aim to perform well relative to their competitors, rather than solely minimizing their completion time.
We first demonstrate that even in basic setups, such as two identical machines with unit-length jobs, a pure Nash equilibrium (NE) assignment may not exist. This observation motivates our inquiry into the complexity of determining whether a given game instance admits a NE. We prove that this problem is NP-complete, even in highly restricted cases. In contrast, we identify specific classes of games where a NE is guaranteed to exist, or where the decision problem can be resolved in polynomial time. Additionally, we examine how competition impacts the efficiency of Nash equilibria, or sink equilibria if a NE does not exist. We derive tight bounds on the price of anarchy, and show that competition may either enhance or degrade overall performance.
@@ 1.3 Related Work
Scheduling games were initially studied in the setting in which each machine processes its jobs in parallel so that the completion time of each job depends on the total load on the machine. The corresponding papers analyze the inefficiency of selfish behavior with respect to the makespan. Czumaj and Vöcking [11] gave tight bounds on the price of anarchy for related machines, whereas Awerbuch et al. [3] and Gairing et al. [15] provided tight bounds for restricted machine settings. These tight bounds grow with the number of machines and motivated the study of coordination mechanisms, i.e., local scheduling policies, to reduce the price of anarchy.
@@ 3.2 Related Machines
    $L_{1}(s)=L_{2}(s)/r$, and each of the two last jobs is prioritized on its machine over the other last job. Formally, $j_{1}\prec_{1}j_{2}$ and $j_{2}\prec_{2}j_{1}$.
@@
    $L_{1}(s)>L_{2}(s)/r$, and either no job on $M_{1}$ has completion time $C_{j_{2}}(s)$, or for the job $j^{\prime}$ such that $C_{j^{\prime}}(s)=C_{j_{2}}(s)$ it holds that $j^{\prime}\prec_{1}j_{2}$.
Proof:    We start by showing that each of these conditions ensures no job satisfies the conditions outlined in Claim 3.1. First, if the slow machine is more loaded, then clearly, $C_{j_{2}}(s)$ is not equal to the completion time of any job on $M_{1}$. Also, $C_{j_{1}}(s)$ is during the interval in which $j_{2}$ is processed on $M_{2}$. Second, if $L_{1}(s)=L_{2}(s)/r$ and $j_{1}\prec_{1}j_{2}$ and $j_{2}\prec_{2}j_{1}$, then none of the two jobs has a beneficial deviation, and no other job is last on its machine. Finally, if the third condition is valid, then as required, no last job has a same-rank job on the other machine, such that a deviation is beneficial.
@@ 5 Related Machines - Equilibrium Existence and Computation
If $0<c<\left\lceil\frac{b}{a}\right\rceil$, assume that Algorithm 2 is executed. After the first $\ell\cdot(a+b)-2$ jobs are assigned, we enter the $({\ell\cdot(a+b)-1})^{th}$ iteration. Let $j_{1}$ be the first unassigned job in $\pi_{1}$. In Step $3$ of the algorithm, machine $i^{*}$ might be either $M_{1}$ or $M_{2}$. Suppose the tiebreaker resolves in favor of assigning $j_{1}$ to $M_{1}$. Let $s$ be the produced schedule, and let $j_{2}$ be the last job processed on $M_{2}$. Since $0<c<\left\lceil\frac{b}{a}\right\rceil$, it follows that $L_{1}(s)>L_{2}(s)/r$. Note that $C_{j_{1}}(s)=C_{j_{2}}(s)$ and $j_{1}\prec_{1}j_{2}$. Therefore, by Theorem 3.3 $s$ is a NE.
@@ 5.1 Global Priority List
For the other direction, let $s$ be a NE schedule, and let $j_{1}$ and $j_{2}$ denote the last jobs processed on $M_{1}$ and $M_{2}$, respectively. Assume by contradiction that $c=0$. Schedule $s$ must be balanced, since otherwise, the last job can benefit from migration. Therefore, $L_{1}(s)=L_{2}(s)/r$, and by Theorem 3.3, since $\pi$ is global, $s$ is not a NE.